# A common first integral from three-body secular theory and Kepler billiards

### v1

@@ 2.5. The Spherical Case
We now consider the system on the sphere $\mathbb{S}^{d}$. The Kepler problem on the sphere has the nice and highly specific property that all non-singular orbits are closed. This allows us to average along all non-singular orbits of the spherical Kepler problem.
@@
By (3), the term $V_{sph,2}$ is linear in $m_{2}$. Note that by analyticity, all the terms and this relationship are well-defined on the whole sphere outside the four singular points of the individual spherical Kepler potentials, even though we have only deduced this formula in a chart for a hemisphere (see relative discussions in [22]). This argument applies also to what follows.
@@
(7) $\{E_{sKep}+V_{sph,2},D+K\}=0.$
The subset in $T^{*}\mathbb{S}^{d}$ such that the spherical Kepler orbits are all non-singular is a dense open set in $T^{*}\mathbb{S}^{d}$. We denote by $\mathcal{I}_{Sph}$ by the set of orbits of $E_{sKep}$ in $T^{*}\mathbb{S}^{d}$ whose projections to $\mathbb{S}^{d}$ pass through other Kepler centers. We set $\mathcal{R}_{sph}=T^{*}\mathbb{S}^{d}\setminus\mathcal{I}_{Sph}$.
@@ 3. Averaging in the Lagrange Problem
We now consider the Lagrange problem in $\mathbb{R}^{2}$. The result and argument extends easily to other cases as well. In $\mathbb{R}^{2}$, two Kepler centers are put at $(0,a)$ and $(0,-a)$ in $\mathbb{R}^{2}$, with masses $m_{1}$ and $m_{2}$ respectively. In addition, a Hooke center is put at $(0,0)$, with a Hooke factor $f$. The metric on $\mathbb{R}^{2}$ is $\|\cdot\|_{a}$, defined as $\|(u,v)\|_{a}=\sqrt{u^{2}+\frac{v^{2}}{1+a^{2}}}$.
@@
$\tilde{V}_{Eucl,2}:=-\dfrac{m_{2}}{\sqrt{x^{2}+\frac{(y+a)^{2}}{1+a^{2}}}}+f(x^{2}+\frac{y^{2}}{1+a^{2}}).$
@@
As in Def. 2.1 we now define
@@
[…]
in which the integration is made with respect to the closed orbits of $E_{Kep}$.
@@
We follow the same proof as that of Prop. 2.2, but now proceed with single frequency averaging along closed Keplerian orbits twice. The first eliminate oscillating part of $m_{2}$-dependent terms, and the second eliminate oscillating part of $f$-dependent terms. We obtain as the proof of Prop. 2.2
$\{\tilde{E}_{Eucl,a},D\}+O(m_{2},f)=0$
and thus
@@ 4. The Systems in Hyperbolic Spaces
As in Killing [13] it is possible to define Kepler, two-center and Lagrange problems in the hyperbolic spaces in a similar fashion. Projective dynamics extends to this case as well. One way to realize the hyperbolic space is to write it as a sheet of the unit pseudosphere in $\mathbb{R}^{d,1}$. The Euclidean-hyperbolic correspondence is now again given via central projection from the origin in $\mathbb{R}^{d,1}$. Analogous of Kepler and Hooke potentials can be defined analogously as in the spherical case, by replacing cotangent and tangent by hyperbolic cotangent and hyperbolic tangent respectively. This way we obtain the same results also for partially averaged systems from two-center problem and Lagrange problem in a hyperbolic space: $D$ is always a first integral of the partially-averaged systems analogously defined. Here the partially-averaged system is again defined in the region such that the Keplerian orbits are closed.
@@ 5.1. Integrable Kepler billiards
Let $(M,g,V)$ be a natural mechanical system, in which $(M,g)$ is a Riemannian manifold and $V:M\mapsto\mathbb{R}$ is the potential. Adding a codimension-1 submanifold $\mathcal{B}$ of $M$ to the data defines a system of mechanical billiards. In such a system, the particle moves under the influence of the potential $V$ and in addition reflected elastically while reaching $\mathcal{B}$. This defines a continuous, non-smooth dynamical systems. The total energy of the system is a preserved quantity of this system. As in the smooth case, the system is called integrable if it possesses $d:=\dim{M}$ preserved quantities which are independent and are in involution with respect to the standard Poisson structure on the cotangent bundle. In particular, if $\dim{M}=2$, then the system is integrable provided a conserved quantity in addition to the energy exists.
@@
We consider a family of confocal conic sections in the plane and we put the Kepler potential at one of their common foci. This always gives rise to integrable Kepler billiards in the plane, with $D$ the additional first integral, in which $h$ is interpreted as the focus-center distance. [22]. Likely these are all the integrable cases [3].
@@
5.2. A billiard system with the partially-averaged system
We shall generalize the notion of integrable mechanical billiard slightly. The underlying mechanical system is now replaced by a smooth, autonomous Hamiltonian system defined on $T^{*}M$. If the configuration space $M$ has dimension 2, such a system is called integrable if it has two independent, commuting conserved quantities.
We now define a set of billiard systems:
In the Euclidean space, on the sphere and in the hyperbolic space, consider any finite combination of open subsets from the family of quadrics defined as in Thm. 5.2. Then, any finite combination of open subsets from this family as reflection wall, with the partially-averaged system as the underlying dynamical system, defines a billiard system. This system mixes the continuous secular evolution of orbits described by the partially-regularized systems with drastic changes of orbits due to reflections with the reflection wall. Such a system does not fit into the above definition of mechanical billiards though, for it is not necessarily defined with a natural mechanical system, whose total energy is the sum of kinetic and potential energies. Nevertheless it still defines a systems combining mechanical and billiard dynamics.
@@ 5.2. A billiard system with the partially-averaged system
For example, this asserts that the system in the plane, in which a particle moves according to the flow of $E_{Eucl,a}$ until it hits a straight line and gets reflected at this line has two conserved quantities, namely the Kepler energy $E_{Kep}$ and $D$. The Hamiltonian $E_{Eucl,a}$ of the system though, is not necessarily preserved under the reflection.
@@
Acknowledgements: L. Z. is supported by the DFG Heisenberg Programme ZH 605/4-1.

### v2

@@ 2.5. The Spherical Case
We now consider the system on the sphere $\mathbb{S}^{d}$. The Kepler problem on the sphere has the nice and highly specific property that all non-singular orbits are closed. Indeed these are all spherical ellipses. This allows us to average along all non-singular orbits of the spherical Kepler problem.
@@
By (3), the term $V_{sph,2}$ is linear in $m_{2}$. Note that by analyticity, all the terms and this relationship are well-defined on the whole sphere outside the four singular points of the individual spherical Kepler potentials, even though we have only deduced this formula in a chart for a hemisphere (see related discussions in [26]). This argument applies as well to what follows.
@@
(7) $\{E_{sKep}+V_{sph,2},D+K\}=0,$
in which the bracket is the Poisson bracket on $T\mathbb{S}^{2}$ transported from $T^{*}\mathbb{S}^{2}$ via Legendre transformation. Note that the same is true for $\mathbb{S}^{d}$ for any $d=1,2,3,\cdots$.
The subset in $T\mathbb{S}^{d}$ such that the spherical Kepler orbits are all non-singular is a dense open set in $T\mathbb{S}^{d}$. We denote by $\mathcal{I}_{Sph}$ by the set of orbits of $E_{sKep}$ in $T\mathbb{S}^{d}$ whose projections to $\mathbb{S}^{d}$ pass through other Kepler centers. We set $\mathcal{R}_{sph}=T\mathbb{S}^{d}\setminus\mathcal{I}_{Sph}$.
@@
Note that on the sphere, the existence of abstract canonical coordinates can be shown in the same way as Lem 2.1, in which no Euclidean assumption has been made.
@@ 3. Averaging in the Lagrange Problem
We now consider the Lagrange problem in $\mathbb{R}^{2}$. The result and argument extends easily to other cases as well. In $\mathbb{R}^{2}$, two Kepler centers are put at $(0,a)$ and (0,a)(0,-a) in $\mathbb{R}^{2}$, with masses $m_{1}$ and $m_{2}$ respectively. In addition, a Hooke center is put at $(0,0)$, with a Hooke factor $f$. The metric on $\mathbb{R}^{2}$ is $\|\cdot\|_{a}$, defined as $\|(u,v)\|_{a}=\sqrt{u^{2}+\frac{v^{2}}{1+a^{2}}}$.
@@
$\tilde{V}_{Eucl,2}:=-\dfrac{m_{2}}{\sqrt{x^{2}+\frac{(y+a)^{2}}{1+a^{2}}}}+f\Bigl(x^{2}+\frac{y^{2}}{1+a^{2}}\Bigr).$
@@
We assume as before $m_{1}>0$, so that $E_{Kep}$ has non-trivial (closed) orbits of negative energy. As in Def. 2.1 we now define
@@
$\tilde{E}_{Eucl,a}=E_{Kep}+\int_{S^{1}}\tilde{V}_{Eucl,2}\,ds,$
in which the averaging is made with respect to the closed orbits of $E_{Kep}$.
@@
We follow the same proof as that of Prop. 2.2, but now proceed with single frequency averaging along closed Keplerian orbits twice. The first eliminate oscillating part of $m_{2}$-dependent terms and the second eliminate oscillating part of $f$-dependent terms. We obtain as the proof of Prop. 2.2
$\{\tilde{E}_{Eucl,a},D\}+O(m_{2},f)\dot{=}0.$
As in previous case, by Lem 2.2 the contributions of order $m_{2}$ and of order $f$ from the $O(m_{2},f)$-term has zero $\tilde{\ell}_{1}$-average, while $\{\tilde{E}_{Eucl,a},D\}$ is independent of $\tilde{\ell}_{1}$. So we conclude as before that
@@ 4. The Systems in Hyperbolic Spaces
As in Killing [15] it is possible to define Kepler, two-center and Lagrange problems in the hyperbolic spaces in a similar fashion. The approach based on projective dynamics extends to this case as well. One way to realize the hyperbolic space is to write it as a sheet of the unit pseudosphere in $\mathbb{R}^{d,1}$. The Euclidean-hyperbolic correspondence is now again given via central projection from the origin in $\mathbb{R}^{d,1}$. Analogous of Kepler and Hooke potentials can be defined analogously as in the spherical case, by replacing cotangent and tangent by hyperbolic cotangent and hyperbolic tangent respectively. This way we obtain the same results also for partially averaged systems from two-center problem and Lagrange problem in a hyperbolic space: $D$ is always a first integral of the partially-averaged systems analogously defined. Here the partially-averaged system is again defined in the region such that the Keplerian orbits are closed.
@@ 5.1. Integrable Kepler billiards
Let $(M,g,V)$ be a natural mechanical system, in which $(M,g)$ is a Riemannian manifold and $V:M\mapsto\mathbb{R}$ is the potential. Adding a codimension-1 submanifold $\mathcal{B}$ of $M$ to the data defines a system of mechanical billiards. In such a system, the particle moves under the influence of the potential $V$ and in addition reflected elastically while reaching $\mathcal{B}$. This defines a continuous, non-smooth dynamical systems. The total energy of the system is a preserved quantity of this system. As in the smooth case, the system is called integrable if it possesses $d:=\dim{M}$ preserved quantities which are independent and are in involution. In particular, if $\dim{M}=2$, then the system is integrable provided a conserved quantity in addition to the energy exists.
@@
Likely these are all the integrable cases with the presence of at least one effective Kepler center [4, 5].
In higher dimensions we have
@@
5.2. A set of billiard systems with the partially-averaged system
We shall generalize the notion of integrable mechanical billiard slightly. The underlying mechanical system is now replaced by a smooth, autonomous Hamiltonian system defined on $TM$ (as previously equipped with a symplectic form from $T^{*}M$ via Legendre transformation). If the configuration space $M$ has dimension 2, such a system is called integrable if it has two independent, commuting conserved quantities.
We now define a set of billiard systems in this sense:
In the Euclidean space, on the sphere and in the hyperbolic space, consider any finite combination of open subsets from the family of quadrics defined as in Thm. 5.2. Then, any finite combination of open subsets from this family as reflection wall, with the partially-averaged system as the underlying dynamical system, defines a billiard system. This system mixes the continuous secular evolution of orbits described by the partially-regularized systems with drastic changes of orbits due to reflections with the reflection wall.
@@ 5.2. A set of billiard systems with the partially-averaged system
For example, this asserts that the system in the plane, in which a particle moves according to the flow of $E_{Eucl,a}$ until it hits a straight line and gets reflected at this line has two conserved quantities, namely the Kepler energy $E_{Kep}$ and $D$. This system combines the work of [19] and [11]. The Hamiltonian $E_{Eucl,a}$ of the system though, is not necessarily preserved under the reflections at the straight line wall.
@@
Acknowledgements: G. P. acknowledges the ERC project 677793 (2016–2022). L. Z. is partially supported by the DFG Heisenberg Programme ZH 605/4-1 and the Fundamental Research Funds for the Central Universities of China.